A 20-g wooden ball is set in horizontal motion using the same set-up used in the activity. The hanging mass is 100 g. The radius of the motion is supposed to be 60 cm however the string made a 100 angle above the horizontal during the actual motion. What is the magnitude of the velocity of the wooden ball?


1
Expert's answer
2021-04-06T13:53:53-0400
"F_T=F_c,""m_hg=\dfrac{m_bv^2}{rcos\theta},""v=\sqrt{\dfrac{m_hgrcos\theta}{m_b}},""v=\sqrt{\dfrac{0.1\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot0.6\ m\cdot cos10^{\circ}}{0.02\ kg}}=5.38\ \dfrac{m}{s}."